Earth is revolving around the sun is the distance of the earth from the sun is reduced one fourth of the present distance than the present in length reduced by

Grade:11

1 Answers

Arun
6 years ago
Dear Hartvik

The ratio of the time periods would be {new is say T1, actual is say T2}

 

 

 

Note that T is directly proportional to √(R3)

 

 

 

T1/T2 = √([R/4]3)/√(R3) = 1/8

 

 

 

Hence T1 = T2/8 = 365/8

 

 = 45.625

 

Which is approx 46 days.

 

Reagrds

Arun(askIITians forum expert)